# Burnside type results for fusion categories

### v10

@@ 5. Kernels of fusion rings, Brauer’s theorem
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let ${(H,\mathcal{B})}_{+}$ be the set of all elements $x\in H$ such that $x=\sum_{b\in{\mathcal{B}}}x_{b}b$ with $x_{b}\in\mathbb{R}_{\geq 0}$. A standard element $b\in\mathcal{B}$ is called a constituent of $x\in{(H,\mathcal{B})}_{+}$ if $x_{b}>0$.
Let $\psi:H\rightarrow\mathbb{C}$ be a linear character, and $E=E_{\psi}\in H$ be its primitive central idempotent. Then, since $E$ is an eigenvector for any $L_{x}$ with $x\in H_{+}$, it follows that $|\psi(x)|\leq\mathrm{FPdim}(x)$, for any object $x\in{(H,\mathcal{B})}_{+}$.
@@
Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset closed under the involution of ${(H,\mathcal{B})}$, and such that $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$. Then ${(L,\;\mathcal{S})}$ is called a sub-hypergroup of ${(H,\mathcal{B})}$.
We mention that the notion of kernels of objects of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
@@
Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$.
@@
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\mid\psi(x)\psi(y)\mid=\mid\psi(xy)\mid=\mid\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\mid$
$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mid\psi(z)\mid\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y),$
It is easy to see that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
Dually, one can define the following:
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. For any $x\in H_{+}$, one can define ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ as the set of all standard elements $\psi\in{\widehat{\mathcal{B}}}$ such that $\psi(x)=\mathrm{FPdim}(x)$.
The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$ if ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
@@ 5.1. Brauer’s theorem for ARN-hypergroup
For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function on $H$ such that $m(\sum_{i=1}^{m}\alpha_{i}x_{i},\sum_{i=1}^{m}\beta_{i}x_{i})=\sum_{i=1}^{m}\frac{\alpha_{i}\beta_{i}}{h_{i}}$. In the case of an abelian hypergroup note that Equation (2.2) implies the following:
(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*})\;\;\text{for all}\;\;x,y\in H.$
@@
${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
@@
Note that $\mathrm{FPdim}(x)=\sum_{\{s\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. If $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
$\mathrm{FPdim}(x)=|\mu_{j}(x)|\leq\sum_{\{s|\;p_{s}>0\}}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x),$
and this implies that $\mu_{j}\in\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. The converse is obvious. ∎
Next theorem is a generalization of Brauer’s theorem. Although the proof is the same as the one from [Bur16], we include it here for the completeness of the paper.
For an element $x\in{(H,\mathcal{B})}_{+}$, we denote by $<x>$ the sub-hypergroup generated by $x$. By definition, it is the sub-hypergroup generated by the constituents of all powers $x^{n}$ with $n\geq 1$.
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $x\in{(H,\mathcal{B})}_{+}$. Then $<x>=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
@@
Suppose that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\mu_{1}$. Write $x=\sum_{j=1}^{m}\mu_{j}(x)F_{j}$ and then
$x^{n}=\sum_{j=1}^{m}\mu_{j}(x)^{n}F_{j}$
Partition the set $\mathcal{J}=\mathcal{J}_{1}\sqcup\mathcal{J}_{2}\sqcup\cdots\sqcup\mathcal{J}_{s}$ such that the value $\mu_{j}(x)$ is constant on each component $\mathcal{J}_{l}$ of the partition. Denote by $\alpha_{l}:=\mu_{j}(x)$ this constant value of $\mu_{j}(x)$ with $j\in\mathcal{J}_{l}$. Note that by the hypothesis, the partition component containing $\{1\}$ is a singleton.
Also note that for any $s\in\mathcal{B}$ and any $a\in H$, then $m(s,a)=\tau(s^{*}a)$. Thus
${m(s,x^{n})}{}=\tau(s^{*}x^{n})=\sum_{j=1}^{m}\mu_{j}(s^{*})\mu_{j}(x)^{n}\tau(F_{j})=\sum_{l\in\mathcal{P}}\big{(}\sum_{j\in\mathcal{J}_{l}}\tau(F_{j})\mu_{j}(s^{*})\big{)}\alpha_{l}^{n}$
Denote
$a_{l}:=\sum_{j\in\mathcal{A}_{l}}\tau(F_{j})\mu_{j}(s^{*}).$
Note that $a_{1}=n{(H,\mathcal{B})}>0$, by Equation (2.9). If $m(s,x^{n})=0$ for all $n$, then
$\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}=0$
for all $n\in\mathbb{N}$, and Vandermonde determinant gives $a_{l}=1$. This is a contradiction since $a_{1}\neq 0$.
For the converse, suppose that $<x>=H$. If $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$, then by Lemma 5.3, $\mu_{j}\in\ker(b)$, for any $b\in\mathcal{B}$. Thus $\mu_{j}=\mu_{1}$. ∎
5.2. The cointegral $\lambda_{(L,S)}$ in terms of kernels
Recall from §1 that for any RN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent corresponding to $\mu_{1}=\mathrm{FPdim}$.
@@ 5.2. The cointegral $\lambda_{(L,S)}$ in terms of kernels
For any $x\in{(H,\mathcal{B})}_{+}$, define $J_{x}:={\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
$\lambda_{<x>}=\sum_{j\in J_{x}}F_{j}$
@@
Suppose that $<x>={(L,\;\mathcal{S})}$. Let $G_{1},\dots,G_{s}$ be the primitive idempotents of ${(L,\;\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters. Moreover, as above one may suppose that $G_{1}=\lambda_{{(L,\;\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
We look at the restrictions of the characters $\mu_{j}$ at ${(L,\;\mathcal{S})}$. Then, there is a surjective function $f:\{1,\dots,r\}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\;\mathcal{S})}}=\psi_{f(j)}$, for all $1\leq j\leq s$.
With the above notation, it will be shown that $f^{-1}(0)=\mathcal{J}_{x}$ Indeed, since $x$ is a faithful element of ${(L,\;\mathcal{S})}$, it follows that $\ker_{{}_{(L,\;\mathcal{S})}}(x)=\{\psi_{1}\}$. On the other hand, note that $\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\;\mathcal{S})}}(x)=\psi_{f(j)}(x)$. Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\;\mathcal{S})}}(x)$, i.e. $f(j)=0$. ∎
@@ 6.1. Notion of the center of a standard element
If ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup, then as in Lemma 5.3, it is easy to see that $Z_{{}_{(H,\mathcal{B})}}(x_{i})$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$.
@@
Again, as in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$ if ${(H,\mathcal{B})}$ is an ARN-hypergroup..
@@ 6.2. Adjoint sub-hypergroup
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}\in H$. Then
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
@@
$|\psi(I(1))|\leq\sum_{i=1}^{m}|h_{i}\psi(x_{i}x_{i^{*}})|=\sum_{i=1}^{m}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
@@
Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $<I(1)>$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
6.3. On the support $J_{{(L,\;\mathcal{S})}}$
@@
Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset $J_{{(L,\;\mathcal{S})}}$ such that
(6.1) ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}=\sum\limits_{j\in{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}}F_{j}.$
The next result is a generalization of [Bur20, Lemma 6.4].
@@
${J_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
@@
By Theorem 5.7, it is enough to show the following:
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}({(H,\mathcal{B})})\}.$
Equation (2.5) gives that $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}$, and therefore, $n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}$ if and only if $|\mu_{j}(x_{i})|^{2}=d_{i}^{2}$, for all $i\in\mathcal{I}$. This is equivalent to $\mu_{j}\in Z_{{}_{(H,\mathcal{B})}}(x_{i})$, for all $i\in\mathcal{I}$. The result follows by Proposition 6.1. ∎
@@
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in J_{{{(H,\mathcal{B})}_{ad}}}$.
@@
It follows from Proposition 6.4 and Lemma 2.24.
@@
(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}\widehat{F}_{j}$

### v11

@@ 5. Kernels of fusion rings, Brauer’s theorem
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let ${(H,\mathcal{B})}_{+}$ be the set of all elements $x\in H$ such that $x=\sum_{b\in{\mathcal{B}}}x_{b}b$ with $x_{b}\in\mathbb{R}_{\geq 0}$. A standard element $b\in\mathcal{B}$ is called a constituent of $x\in{(H,\mathcal{B})}_{+}$ if $x_{b}>0$. Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$. Let $\psi\in{\widehat{\mathcal{B}}}$. For all $x\in{(H,\mathcal{B})}_{+}$,
$|\psi(x)|=|\psi(\sum_{b\in{\mathcal{B}}}x_{b}b)|\leq\sum_{b\in{\mathcal{B}}}x_{b}|\psi(b)|\leq\sum_{b\in{\mathcal{B}}}x_{b}\mathrm{FPdim}(b)=\mathrm{FPdim}(x),$
by Frobenius-Perron theory.
@@
Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset such that
  • $\mathcal{S}$ is closed under the involution of ${(H,\mathcal{B})}$,
  • $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$.
Then ${(L,\mathcal{S})}$ is called a sub-hypergroup of ${(H,\mathcal{B})}$.
The notion of kernel of object of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
@@
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\left|\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\right|$
$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}|\psi(z)|\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y).$
It is easy to deduce that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
Dually, we can define the following:
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let $x\in H_{+}$. Define
${\ker_{{}_{{(H,\mathcal{B})}}}}(x):=\{\psi\in{\widehat{\mathcal{B}}}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of $^{\widehat{(H,\mathcal{B})}}$ if $^{\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
@@ 5.1. Brauer’s theorem for ARN-hypergroup
For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function $m$ on $H$ by
$m(\sum_{i\in\mathcal{I}}\alpha_{i}x_{i},\sum_{i\in\mathcal{I}}\beta_{i}x_{i})=\sum_{i\in\mathcal{I}}\frac{\alpha_{i}\beta_{i}}{h_{i}}.$
In the abelian case, Equation (2.2) implies
(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*}),\;\;\text{for all}\;\;x,y\in H.$
@@
${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s\in\mathcal{B}|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
@@
Note that $\mathrm{FPdim}(x)=\sum_{\{s\in\mathcal{B}\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. Let $\mu_{j}\in\widehat{B}$ then
$|\mu_{j}(x)|\leq\sum_{\{s\in\mathcal{B}|\;p_{s}>0\}}p_{s}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x).$
Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\mu_{j}\in\bigcap_{{}_{\{s\in\mathcal{B}|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. ∎
The following theorem is a generalization of Brauer’s theorem. Although the proof is the same as that in [Bur16], we include it here for the sake of completeness.
For $x\in{(H,\mathcal{B})}_{+}$, we denote by $\langle x\rangle$ the sub-hypergroup generated by $x$, i.e. whose basis is the set of constituents of all powers $x^{n}$, with $n\geq 1$.
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup, and $x\in{(H,\mathcal{B})}_{+}$. Then $\langle x\rangle=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
@@
Partition the set $\mathcal{I}=\bigsqcup_{l\in\mathcal{P}}\mathcal{I}_{l}$ such that the values $\mu_{j}(x)$ are constant on each component $\mathcal{I}_{l}$ of the partition, but distinct otherwise (i.e. for all $l,l^{\prime}\in\mathcal{P}$, $j\in\mathcal{I}_{l}$, $j^{\prime}\in\mathcal{I}_{l^{\prime}}$, $\mu_{j}(x)=\mu_{j^{\prime}}(x)$ if and only if $l=l^{\prime}$). Let $s\in\mathcal{B}$. By Equation (5.1),
${m(x^{n},s)}{}=\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x)^{n}\mu_{j}(s^{*})=\sum_{l\in\mathcal{P}}\left(\sum_{j\in\mathcal{I}_{l}}\frac{1}{n_{j}}\mu_{j}(s^{*})\right)\alpha_{l}^{n}=\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}$
with
$a_{l}:=\sum_{j\in\mathcal{I}_{l}}\frac{1}{n_{j}}\mu_{j}(s^{*}),\text{ and }\alpha_{l}:=\mu_{j}(x)\text{ when }j\in\mathcal{I}_{l}.$
Note that $s\not\in\langle x\rangle$ if and only if $m(x^{n},s)=0$ for all $n$, if and only if $Mv=0$, with $M$ be the matrix $(\alpha_{l}^{n})$ and $v$ be the vector $(a_{l})$. But $M$ is invertible because its determinant is nonzero, because all $\alpha_{l}$ are distinct (by Vandermonde determinant formula). Thus, $s\in\langle x\rangle$ if and only if there is $l\in\mathcal{P}$ such that $a_{l}$ is nonzero.
[…]
Finally, $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$ if and only if $x\in\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})$, if and only if $\langle x\rangle\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})$ by Lemma 5.3. So if $\langle x\rangle=H$ then $\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})=H$, i.e $\mu_{j}=\mu_{1}$. ∎
5.2. The cointegral $\lambda_{(L,\mathcal{S})}$ in terms of kernels
Recall from §1 that for any ARN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent $F_{1}$ corresponding to $\mu_{1}=\mathrm{FPdim}$. Let ${(L,\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$, and let $\lambda_{(L,\mathcal{S})}$ be the idempotent in $L\subset H$ corresponding to $\mu_{1}\big{|}_{(L,\mathcal{S})}$. By Equation (2.4), we get that:
(5.2) $\lambda_{(L,\mathcal{S})}=\frac{1}{n{(L,\mathcal{S})}}\sum_{x_{i}\in\mathcal{S}}h_{i}d_{i}x_{i}.$
In particular, $\lambda_{(L,\mathcal{S})}=\lambda_{(K,\mathcal{T})}$ if and only if ${(L,\mathcal{S})}={(K,\mathcal{T})}$.
@@ 5.2. The cointegral $\lambda_{(L,\mathcal{S})}$ in terms of kernels
For any $x\in{(H,\mathcal{B})}_{+}$, let $\mathcal{I}_{x}$ be the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$. Then
$\lambda_{\langle x\rangle}=\sum_{j\in\mathcal{I}_{x}}F_{j}$
@@
Suppose that $\langle x\rangle={(L,\mathcal{S})}$. Let $G_{1},\dots,G_{s}$ be the primitive central idempotents of ${(L,\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters. As above, let us take $G_{1}=\lambda_{{(L,\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
We look at the restrictions of the characters $\mu_{j}$ at ${(L,\mathcal{S})}$. Then, there is a surjective function $f:\mathcal{I}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\mathcal{S})}}=\psi_{f(j)}$, for all $j\in\mathcal{I}$.
With the above notation, we are reduced to show that $f^{-1}(\{1\})=\mathcal{I}_{x}$ By Theorem 5.6, since $\langle x\rangle={(L,\mathcal{S})}$, then $\ker_{{}_{(L,\mathcal{S})}}(x)=\{\psi_{1}\}$. On the other hand,
$\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\mathcal{S})}}(x)=\psi_{f(j)}(x).$
Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\mathcal{S})}}(x)$, i.e. $f(j)=1$. ∎
@@ 6.1. Notion of the center of a standard element
As in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$. Idem for $Z_{{}_{(H,\mathcal{B})}}(x_{i})$, if $^{\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
@@ 6.2. Adjoint sub-hypergroup
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}\in H$. Then
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i\in\mathcal{I}}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
@@
$\psi(I(1))=\sum_{i\in\mathcal{I}}h_{i}\psi(x_{i}x_{i^{*}})=\sum_{i\in\mathcal{I}}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i\in\mathcal{I}}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i\in\mathcal{I}}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
@@
Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $\langle I(1)\rangle$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
6.3. On the support ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$
@@
Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}\subset\mathcal{I}$ such that
(6.1) ${\lambda_{{}_{{(L,\mathcal{S})}}}}=\sum_{j\in{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}}F_{j}.$
In particular, by Equation (5.2), ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}={\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}$ if and only if ${(L,\mathcal{S})}={(K,\mathcal{T})}$.
The next result generalizes [Bur20, Lemma 6.4].
@@
${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{I}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
@@
By Theorem 5.7 and Definition 6.2, ${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}={\mathcal{I}}_{I(1)}$, the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Thus,
${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{I}\;|\;\mu_{j}(I(1))=\mathrm{FPdim}(I(1))\}.$
But $\mathrm{FPdim}(I(1))=\mathrm{FPdim}{(H,\mathcal{B})}$, and $\mu_{j}(I(1))=n_{j}$ by Equation (2.5). ∎
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in{{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}$.
@@
It follows from Proposition 6.4 and Lemma 2.31. ∎
@@
(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}$